# Exciton Condensation in Landau Levels of Quantum Spin Hall Insulators

### v1

@@ Abstract
Adding this term to the mean-field Hamiltonian, the self-consistency equation of the exciton condensation is modified as follows,
$\Delta=\frac{A+\lambda}{2E^{\prime}}\big{(}f(-E^{\prime})-f(E^{\prime})\big{)},$ (18)
in which the modified quasiparticle excitation energy $E^{\prime}=\frac{1}{2}\sqrt{(\epsilon_{0}^{+}-\epsilon_{0}^{-})^{2}+4|A+\lambda|^{2}}$. As shown in Fig. 2 (c), the exciton condensation order $\Delta$ is nonzero at any temperature and magnetic field due to the explicit symmetry breaking, thus the EI phase boundary is smeared. Nonetheless, the exciton condensation order is strong only in the vicinity of $B_{\mathrm{c}}$ at low temperature and is vanishingly small elsewhere, thus the feature of the ESR spectroscopy is not significantly affected by the weak inversion asymmetry.
Conclusion.— In summary, we theoretically studied the 2D QSHI with the space inversion and the uniaxial spin rotation symmetries in a perpendicular magnetic field. In the noninteracting case, the system undergoes a topological transition from the QSHI to the NI state at a critical magnetic field $B_{\mathrm{c}}$. The Coulomb interaction triggers the exciton condensation in the vicinity of $B_{\mathrm{c}}$ at low temperature and spontaneously breaks the inversion and the spin rotation symmetries. We propose that the ESR spectroscopy in the Voigt configuration can directly probe the exciton condensation order. Our results should apply to the InAs/GaSb QWs and other QSHIs with (approximate) inversion and spin rotation symmetries.
@@ Acknowledgements.
H.-M.P. thanks Jin-Tao Jin for helpful discussions. L.Z. is grateful to Rui-Rui Du, Zhongdong Han and Tingxin Li for related stimulating collaborations, and to Yuan Li and Jia-Wei Mei for helpful discussions. This work is supported by the National Natural Science Foundation of China (No. 12174387), the Chinese Academy of Sciences (Nos. XDB28000000, YSBR-057, and JZHKYPT-2021-08), and the Innovative Program for Quantum Science and Technology (No. 2021ZD0302600). L.Z. is also supported by the Xiaomi Foundation.

### v2

@@ Abstract
Adding this term to the mean-field Hamiltonian, the self-consistency equations of the exciton condensation are modified as follows,
$\Delta_{k}=\frac{A_{k}+\lambda}{2E_{k}^{\prime}}\big{(}f(-E_{k}^{\prime})-f(E_{k}^{\prime})\big{)},\quad\forall k,$ (18)
in which the modified quasiparticle excitation energy $E_{k}^{\prime}=\frac{1}{2}\sqrt{(\epsilon_{0}^{+}-\epsilon_{0}^{-})^{2}+4|A_{k}+\lambda|^{2}}$. The modified equations are also solved by a constant $\Delta_{k}=\Delta$; However, as shown in Fig. 2 (e), $\Delta$ is nonzero at any temperature and magnetic field due to the explicit symmetry breaking, thus the EI phase boundary is smeared. Nonetheless, the exciton condensation order is strong only in the vicinity of $B_{\mathrm{c}}$ at low temperature and is vanishingly small elsewhere, thus the main feature of the ESR spectroscopy is not significantly affected by the weak inversion asymmetry, which is illustrated in Fig. 2 (g,h) 22footnotemark: 2.
Conclusion.— In summary, we theoretically studied the 2D QSHI with the space inversion and the uniaxial spin rotation symmetries in a perpendicular magnetic field. In the noninteracting case, the system undergoes a topological transition from the QSHI to the NI state at a critical magnetic field $B_{\mathrm{c}}$. The Coulomb interaction triggers the exciton condensation in the vicinity of $B_{\mathrm{c}}$ at low temperature and spontaneously breaks the inversion and the spin rotation symmetries. We propose that the ESR spectroscopy with the ac magnetic field also aligned in the perpendicular direction (the Voigt configuration) can directly probe the exciton condensation order. Our results should apply to QSHIs with (approximate) inversion and/or uniaxial spin rotation symmetries such as the InAs/GaSb QWs and the monolayer transition metal chalcogenides.
@@ Acknowledgements.
H.-M.P. thanks Jin-Tao Jin for helpful discussions. L.Z. is grateful to Rui-Rui Du, Zhongdong Han and Tingxin Li for related stimulating collaborations, and to Yuan Li and Jia-Wei Mei for helpful discussions. This work is supported by the National Natural Science Foundation of China (No. 12174387), the Chinese Academy of Sciences (Nos. XDB28000000, YSBR-057, and JZHKYPT-2021-08), the University of Chinese Academy of Sciences, and the Innovative Program for Quantum Science and Technology (No. 2021ZD0302600). L.Z. is also supported by the Xiaomi Foundation.